# Metadata practices for simulation workflows

### v1

@@ Abstract
Computer simulations are an essential pillar of knowledge generation in science. Understanding, reproducing, and exploring the results of simulations relies on tracking and organizing metadata describing numerical experiments. However, the models used to understand real-world systems, and the computational machinery required to simulate them, are typically complex, and produce large amounts of heterogeneous metadata. Here, we present general practices for acquiring and handling metadata that are agnostic to software and hardware, and highly flexible for the user. These consist of two steps: 1) recording and storing raw metadata, and 2) selecting and structuring metadata. As a proof of concept, we develop the Archivist, a Python tool to help with the second step, and use it to apply our practices to distinct high-performance computing use cases from neuroscience and hydrology. Our practices and the Archivist can readily be applied to existing workflows without the need for substantial restructuring. They support sustainable numerical workflows, facilitating reproducibility and data reuse in generic simulation-based research.
1 Introduction
Recent advances in high-performance computing (HPC) technology enable simulations of increasingly large and complex models, which offer huge potential for science and society but are more challenging to accurately reproduce, share, and interpret. These difficulties make user stories such as the following common among researchers:
@@ 1 Introduction
    Scientist X cannot reproduce the results of scientist Y due to a lack of information on software dependencies and inconsistencies between the information provided in the article and in the code published by Y. Even personal communication with Y does not resolve these inconsistencies. [1]
@@
    Each member of a group of scientists regularly runs simulations of the same mathematical model to investigate different scientific questions. Because the data generated by each simulation is similar and could be useful to others, sharing it is desirable to minimize time and energy costs. However, the scientists have no efficient way of communicating the information necessary to understand the structure of each dataset and the details of how it was generated[2].
@@
    A group of scientists is developing simulation software. After each development cycle, the group runs a set of benchmarking experiments with different configurations and models to continuously monitor software performance (“continuous benchmarking”[3]). After years of development, the group has accumulated large amounts of benchmarking data for each software version. While exploring the data for an individual software version is simple, the scientists have no efficient way of comparing versions which use similar configurations or model types.
@@
One would assume that the above challenges should not exist in simulation science, due to our perception of full control over digital implementations[4]. However, this assumption is wrong for two reasons. First, users often are not aware of every aspect of their hardware and software systems, for example low-level hardware settings[5], implementation details of a software package, or the implications of using a specific operating system[6]. These aspects are of particular importance for HPC-enabled simulations, where highly specialized hardware and software solutions and the distribution of code among multiple processors may affect the exact simulation outcome or performance measures. Given the limited lifetime of HPC systems, reproducibility may be practically impaired. Second, users often customize aspects of their system without properly documenting their changes, for example modifying the behavior of off-the-shelf software. Therefore, despite the digital nature of simulation research, it remains difficult to acquire and organize metadata describing the details of hardware systems, software stacks, model descriptions and simulation workflows that are necessary to replicate, understand, and share numerical experiments[7, 8, 1, 9]. These metadata arise at all stages of the simulation workflow, including defining and implementing a model; preparing software; generating and executing a job; post-processing, analyzing, and visualizing data; and organizing and storing data. Successfully capturing comprehensive metadata and provenance information along the processing chain ensures reproducibility and replicability[7, 10, 8, 11, 12], allows assessment of simulation outcomes[13, 1], and helps scientists explore, share, and reuse data[14].
@@
Recently, several powerful tools have been developed to organize, execute, and track complex workflows, including their resulting data and metadata, such as Sumatra[10], AiiDA[15], Snakemake[16], and Datalad[17]. Although these tools are useful for new projects, integrating them into existing workflows requires major restructuring. As a consequence, scientists often write custom code for metadata handling to compromise between adding functionality and keeping an established workflow. The resulting inconsistency in metadata management makes it difficult to transfer knowledge between researchers.
@@
Here, we propose a set of practices for flexible metadata management which can readily be integrated into existing simulation workflows without substantial refactoring. The practices are generic and apply to diverse research fields. They advise researchers about which metadata to collect at each stage of a simulation workflow, how to collect them, and how to process the metadata so that they enrich data. We further describe the Python tool Archivist which helps selecting and structuring metadata. Eventually, we apply the proposed practices and the Archivist to a minimal illustrating example, and to two real-world use cases from neuroscience and hydrology.
2 Metadata management practices
@@
To better understand the metadata collection and handling process, we consider a generic knowledge production workflow divided into three sub-workflows (fig. 1): a simulation experiment containing processes generating data and metadata, metadata post-processing where heterogeneous metadata is processed into meaningful formats and structures, and usage of enriched data. Although the implementation of these processes can and will vary across different use cases, this generic knowledge production workflow encompasses archetypes of the components occurring in many implemented workflows.
@@
Figure 1: Metadata management in a generic knowledge production workflow. We conceptually divide the workflow into three consecutive sub-workflows: simulation, metadata post-processing, and data usage. Simulation collects raw metadata (yellow) at each step (blue), then stores it along with a data model describing the structures of data and raw metadata. Post-processing (red) selects raw metadata, then parses and structures it; annotation links this processed metadata to the data. During post-processing, user-defined configuration specifies which raw metadata to select and how to process it, ensuring the final annotated data is suitable for its application. Some of these steps (light red) can be performed by the post-processing framework presented in “Metadata post-processing framework – the Archivist”. Data usage (green) analyzes data to draw conclusions through presentation and comparison; data may also be shared with others. Ultimately, the results of data usage inform subsequent simulations (green arrow). Each sub-workflow consists of autonomous steps (rectangles), user-configured steps (rounded rectangles), data storage (cylinders), and data transfer (arrows). This diagram aims to represent generic simulation-based research; specific workflows may omit some elements (asterisks).
@@
A generic simulation experiment comprises a sequence of steps (fig. 1, Simulation section). We define these steps by abstracting the processes performed in two domain-specific workflows from neuroscience and hydrology[14, 18]. Depending on the complexity of the preparation required before and after the execution of the simulation there might be additional steps, but at minimum we consider the necessary steps to be configuration of the software environment, simulation engine, and model; execution of the simulation; and storage of the generated data and metadata. For clarity we separate data and metadata storage in fig. 1, however this is an implementation choice left to the discretion of the user and for the remainder of this paper we refer to both storages together as one collection for simplicity.
@@
After the simulation experiment, users can analyze the resulting data and metadata collection to gain knowledge on the simulated model or simulation technology. The methods through which data can be analyzed are tightly related to the objectives and intentions of the user. Furthermore, in some disciplines, these methods may require a complex workflow and perhaps even the use of High-Performance Computing (HPC) resources. Since we cannot exhaustively define the different existing strategies for processing data, we provide an abstract description of how the generated data can be exploited (fig. 1, Data usage section). For instance, data is analyzed in order to measure metrics related to model predictions or accuracy, or simulation performance. Graphical visualizations are generated for easier human interpretation. Additionally, differences between multiple versions of simulators, models or parameters are obtained by comparing with previous results. New insights are shared with peers to disseminate knowledge.
@@
As processing the data may inadvertently remove or add specific features that alter the original information, storing the original data allows independent users to exploit individual datasets with different methods, or to process multiple datasets at once, which increases re-usability of the simulation results. Over time the data and metadata generated by different simulation experiments will be compiled and stored in a continuously growing collection. Establishing a data model is necessary to maintain the organization and integrity of this collection. A data model is an abstract representation of data objects and their relationships that organizes relevant attributes of data and standardizes how they relate to each other[19]. This serves as a blueprint for structuring data elements and their relationships, providing a clear understanding of how data is organized, stored, and utilized within a collection. Without the data model, users may not be able to establish an overview of the actual contents and may fail to efficiently explore the collection.
@@
To implement and perform this diversity of data usage tasks metadata is required to properly describe the simulation experiment. When exploring multiple sets of generated data, metadata categorizing each dataset allows researchers to efficiently differentiate them. Moreover, when sharing the results, metadata enhances the knowledge transfer between data users by detailing what, how, and where the experiment was performed. Most of this metadata, can only be found when performing the experiment and can exist at any step of the simulation (fig. 1, blue Data generation boxes). The collection of metadata and their interpretation often varies and evolves with the systems and software used (fig. 1, yellow Raw metadata collection boxes). We describe below some general practices for successfully managing the variety of metadata that arises during simulation.
@@
Metadata are data that describe data. Here, we focus on information that can be directly collected in the environment where the experiment takes place (sometimes termed "hard" metadata [20]). However, the definition of what precise pieces of information should be represented as metadata is difficult to generalize. Each experiment generates an arbitrary amount of information – separating this information into metadata and data depends on the nature and goal of the workflow itself. For instance, if the goal of a simulation workflow is to collect and analyze the data generated by the underlying model, then the performance of the simulation (time to solution, memory consumption, etc.) would be part of the metadata. On the contrary, if the goal of the workflow is to benchmark the simulator then the data generated by the model is of secondary importance, as the recorded performance data would be the desired primary result. Limiting the amount of information collected during a simulation experiment restricts how its output can be used in the future. Therefore, to maximize the potential for reuse of shared experiment results, it is important to collect as much information as possible[10]. Yet, the full extent of this collection needs to be determined on a case by case basis. Sometimes, collecting the full breadth of information available can be redundant. For instance, when simulation models grow complex, the number of available parameters grows in volume too. Well-documented parameter sets often use default values suitable for their corresponding model; during experiments, only a subset of these may actually be modified. In this case, collecting only the modified parameters may be sufficient. Conversely, there may be instances where keeping detailed track of the differences between experiment is necessary. Experience has shown, for example, that differences in software versions may lead to discrepancies in experiment results[1]. In the case of rapidly evolving software such as the model, simulator, or their library dependencies, it is beneficial to keep track of the different versions used for each experiment. As a compromise, we propose collecting information as sustainably as possible to an extent where all currently known user needs are satisfied.
@@
Given the rapidly evolving software and hardware platforms available for simulation alongside user specific set-ups, it is impossible to exhaustively list all collectible metadata. Instead, we categorize the different sources of metadata according to each step of the simulation experiment (Fig. fig. 1, blue Data generation boxes):
@@
  • Software preparation: In this step all the required software (libraries, tools, and programs) are fetched, if needed compiled, and loaded into the environment. Information to collect here includes the source of the software, the parameters related to compilation, and the environment variables defined when loading the software.
  • Job generation (if applicable): In HPC scenarios, job schedulers are generally used to efficiently allocate and share resources of compute clusters. Collectable metadata here include job parameters: machine name, version and partition used, amount and type of resources used, resource configuration such a process binding or thread pinning, and job time limit.
  • Job execution: The main part of the simulation, the execution involves instantiating the simulation model and running the simulation. It is at this step that recording the status of the execution environment itself is the most interesting and least redundant. In contrast to Software preparation, the metadata collected in this step should focus on the environment variables, libraries, and processes present during execution. Additionally, execution time, resource consumption (memory, storage, network, etc.), and system load can only be recorded at this step.
  • Post-processing (if applicable): Depending on the goal and the implementation of the workflow, the data that is generated must be post-processed for further use, which may involve steps such as checking, cleaning, compressing or transforming the data. If post-processing is necessary, documenting the procedures performed on data is essential to have an overview of how the final result was obtained. Thus the information on the different methods used during this step must be recorded.
  • Data persistence: In this step, information on the storing process can be recorded. For example: who generated the data and metadata to be stored? When was it generated and when is it being stored? Where was it generated and where is it being stored to e.g. a file server, a local or remote database? In the case of a file server, the address of the server and the absolute path of the stored files should be recorded. When using a database, new entries are indexed with a unique identifier; storing this identifier allows linking with future related entries and should be recorded correspondingly.
@@
There are two possible ways of collecting metadata: retrieval by a dedicated parallel process “pulling” the relevant pieces of information from each simulation step, or autonomous “pushing” by each component to a metadata storage. The pull-based approach relies on an overarching metadata-monitoring module that is adapted to the specific simulation. The push-based approach, in contrast, does not require such infrastructure at run time. At each step, the governing process simply dumps its metadata in some arbitrary, more or less raw format, independent of other simulation steps. The push-based approach is therefore much more flexible and can readily be applied to existing simulations. With each component presenting its own metadata it is likely that the final collection of metadata contains different structures and formats. Although it is possible to standardize the presentation of metadata across all components, this process would require transformations that are dependent on the objective for which the data is being generated, and can vary from user to user, or interoperability standard that the researcher must adhere to. In some situations, transformations of metadata may even depend on other metadata, making it difficult to perform an ad-hoc transformation process. We therefore propose extracting and structuring the raw metadata in a separate step after the simulation. This presents a clear advantage by decoupling metadata collection processes from user goals and processing software. Collecting as much metadata as possible is essential as it is impossible to get the corresponding metadata after a simulation is performed. Thus, gathering the metadata regardless of format and storing it during simulation for later processing is an efficient way to keep track of the information describing the experiment. We call the collection of these unprocessed metadata files the "raw" metadata (fig. 1, yellow Raw metadata collection boxes).
@@
Transforming the heterogeneous raw metadata into a unified comprehensive format is essential in order to be able to exploit it alongside the data. Given that each workflow is run with a specific objective in mind, the collection of raw metadata contains often much more than needed. Although this is desirable to maximize possible reusability of simulation data, for a given objective the raw metadata should be filtered. For this, a dedicated read operation that extracts the relevant entries from a collection of metadata is performed according to given objective. The extracted metadata is parsed and might be immediately usable depending on the output format of the parsing operations, however a specific structure might be needed to increase interoperability between the processed metadata and other data exploitation software. Therefore, structuring the metadata according to a standard is often a beneficial step to easily and unambiguously leverage the metadata information. With this, the previously heterogeneous metadata collection is now unified, possibly structured according to a standard, and ready to be exploited.
@@
Before being able to use processed metadata and data in unison these need to be combined together. Here we refer to combining as a referencing procedure in which any data entity that makes up a complete data record is cross-referenced with any metadata (item or collection) that assists in using and interpreting the data entity. Although it is possible to exploit metadata and data independently for specific purposes, combining them quickly becomes necessary when the volume and dimension of a dataset starts to grow. When amassing data from multiple simulations, each with varying degrees of different configuration and or models, it becomes apparent that simulation results need to be explored with a higher level of description. For this, "tagging" or annotating the data with the processed metadata creates an enriched dataset. When appropriately stored, users can exploit indexing engines to navigate through these enriched datasets and generate complex data projections. Again, the level of granularity with which data entities are identified and matched to metadata will need to be decided based on the objective of the workflow.
@@
Following these practices will help in handling large heterogeneous volumes of metadata, and combining datasets with processed metadata increases their explorability and exploitability. The actual methods through which these practices can be followed or implemented depend on the technology used to manage the simulation experiment and data processing pipeline. In particular some existing technologies offer support for these practices and are addressed in the “Discussion”.
3 Metadata post-processing framework – the Archivist
@@
In the previous section we showed that, to cope with the heterogeneity of collected metadata, it is necessary to perform parsing and possibly structuring operations to obtain a unified, exploitable metadata representation. Although there exist tools capable of doing these kinds of operations, e.g.,AiiDA[15] parsers (https://.readthedocs.io/projects/-core/en/latest/reference/apidoc/.parsers.html, last access: 23 February 2024), DataLad[17] metadata extractors (https://docs.datalad.org/projects/metalad/en/latest/extractors.html, last access: 23 February 2024), these can pose fixed constraints on the processing of the metadata, may be an inseparable component of a larger framework and may not be used independently.
@@
Here we propose the Archivist as a framework that can be used to perform parsing and structuring operations. It is capable of individually parsing generic files and combining arbitrary parsed results in a customizable structure. We note that data annotation, i.e., linking metadata records to relevant parts of the data in the post-processing pipeline (fig. 1, red Annotation box), is currently not supported because annotations of data need to be heavily customized to fit the underlying structure of the data and this process is difficult to formalize. For example, while data structured according to HDF5-based[21] formats like the NIX[22] standard can make use of intrinsic mechanisms provided for linking collections of metadata entries to data, this mechanism cannot be generalized to other file formats and data models.

### v2

@@ Abstract
Computer simulations are an essential pillar of knowledge generation in science. Exploring, understanding, reproducing, and sharing the results of simulations relies on tracking and organizing the metadata describing the numerical experiments. The models used to understand real-world systems, and the computational machinery required to simulate them, are typically complex, and produce large amounts of heterogeneous metadata. Here, we present general practices for acquiring and handling metadata that are agnostic to software and hardware, and highly flexible for the user. These consist of two steps: 1) recording and storing raw metadata, and 2) selecting and structuring metadata. As a proof of concept, we develop the Archivist, a Python tool to help with the second step, and use it to apply our practices to distinct high-performance computing use cases from neuroscience and hydrology. Our practices and the Archivist can readily be applied to existing workflows without the need for substantial restructuring. They support sustainable numerical workflows, fostering replicability, reproducibility, data exploration, and data sharing in simulation-based research.
Introduction
Recent advances in high-performance computing (HPC) technology enable simulations of increasingly large and complex models. While these simulations offer huge potential for science and society, it becomes more and more challenging to replicate and reproduce the results of such numerical experiments, to efficiently explore the simulation data, and to share them with other scientists.
User stories such as the following are therefore common among researchers:
@@ Introduction
    Replicating results: Scientist X cannot replicate the results of scientist Y due to inconsistencies between the information provided in the scientific article and in the associated code published by Y. Even personal communication with Y does not resolve these inconsistencies[1].
@@
    Data sharing: Each member of a group of scientists regularly runs simulations of the same mathematical model to investigate different scientific questions. The datasets generated by each of these scientists are similar and potentially useful to other members of the group, although what one scientist considers to be metadata can be analyzed as primary data by another scientist with a different objective. It is therefore desirable to share these data to minimize time and energy costs and ultimately to increase scientific productivity. However, the scientists have no efficient way of communicating the information necessary to understand the structure of each dataset[2].
@@
    Data exploration: A group of scientists is developing a simulation software. After each development cycle, the group runs a set of benchmarking experiments with different configurations and models to continuously monitor the simulation software performance (“continuous benchmarking”[3]). After years of development, the group has accumulated large amounts of benchmarking data for each software version. The scientists have no way to easily search and view subsets of the accumulated data, for example, to compare different versions using similar configurations or model types.
@@
One would assume that the above problems, known to different extents in various research fields, should not exist in simulation science, due to our perception of full control over digital implementations[4]. However, this assumption is typically wrong for two reasons. First, users are often not aware of every aspect of their hardware and software systems, such as low-level hardware settings[5], implementation details of a software package, or the implications of using a specific operating system[6]. These aspects are of particular importance for HPC-enabled simulations, where highly specialized hardware and software solutions and the distribution of code among multiple processors may affect the exact simulation outcome or performance measures. Given the limited lifetime of HPC systems, long-term repeatability is practically impaired. Second, users often customize aspects of their system without properly documenting their changes, for example modifying the behavior of off-the-shelf software. Therefore, despite the digital nature of simulation research, it remains difficult to acquire and organize metadata describing the details of hardware systems, software stacks, model descriptions and simulation workflows that are necessary to replicate, reproduce, understand, and share the results of numerical experiments[7, 8, 1, 9]. These metadata arise at all stages of the simulation workflow, including defining and implementing a model; preparing software; generating and executing a job; post-processing, analyzing, and visualizing data; and organizing and storing data.
Throughout this paper, we use the terms ‘reproducibility’ and ‘replicability’ as defined by the Association for Computing Machinery (ACM) (https://www.acm.org/publications/policies/artifact-review-and-badging-current, last access: 12 February 2025). According to these guidelines, results from a computational experiment are reproduced if an independent team obtains the same results using the same experimental setup, whereas results are replicated if an independent team obtains the same results using a different experimental setup. Here, ‘experimental setup’ refers to the artifacts used to generate results, which “can be software systems, scripts used to run experiments, input datasets, raw data collected in the experiment, or scripts used to analyze results”. For simulation experiments, we consider such artifacts to be simulation engines, model and workflow implementations, and their respective configurations. However, it is important to consider not only the software, but also the hardware system used to obtain results. For instance, performance metrics such as time to solution or energy consumption cannot be interpreted without knowledge of the hardware on which the experiment is run. These considerations may blur the conceptual line between reproduction and replication. It is, for example, not always clear whether a small change in the configuration or the hardware leads to a different experimental setup. Ultimately, it is the research question underlying a simulation experiment that determines which aspects of the experimental setup are important to consider when reproducing or replicating results.
Successfully capturing comprehensive metadata and provenance information describing the experiment and its setup is a prerequisite for reproducibility and replicability[7, 10, 8, 11, 12, 9], allows assessment of simulation outcomes[13, 1], and helps scientists explore, share, and reuse data[14].
@@
Recently, several powerful tools have been developed to organize, execute, and track complex workflows, including their resulting data and metadata, such as Sumatra[10], AiiDA[15], Snakemake[16], and DataLad[17]. Although these tools are useful for new projects, using them in existing projects with different, possibly custom, workflow management systems requires deciding whether to rebuild the workflow from scratch using the desired tools or look for alternative solutions.. As a consequence, scientists often write custom code for metadata handling to compromise between adding functionality and keeping an established workflow. The resulting inconsistency in metadata management makes it difficult to transfer knowledge between researchers.
Other community efforts have focused on defining vocabularies or schemas such as RO-Crate[18], CodeMeta (https://codemeta.github.io, last visited: 23 January 2025), or Bioschemas[19]. Such standards enable interoperability of data and metadata produced by computational workflows, and provide practical ways for developers and users to describe and document their experiments. Nonetheless, detailed information on individual workflow executions can only be collected during runtime and often requires additional transformations to comply with vocabularies or schemas. Examples of such information are: the software environment at runtime, computing capabilities of the hardware system in use, or performance and resource consumption of computationally heavy workflow steps. Furthermore, identifying where such information can be found and the methods to transform it depend on the workflow implementation and system used. This dependency makes found solutions harder to transfer to other workflows or systems, and consequently researchers avoid collecting and handling such information unless required by their scientific objectives.
@@
Here, we propose a set of practices for flexible metadata management which can readily be integrated into existing simulation workflows without substantial refactoring. The practices apply to commonly used workflows across diverse research fields. They advise researchers about which metadata to collect at each stage of a simulation workflow, how to collect it, and how to process the metadata so that they enrich data. We further describe our Python tool Archivist which helps process metadata files using user-defined functions and combine the outputs into a unified file. Finally, we apply our proposed practices and the Archivist to a minimal illustrating example, and to two real-world use cases from neuroscience and hydrology.
Results
Metadata management practices
@@ Results
To better understand the metadata collection and handling processes, we consider a generic knowledge production workflow divided into three sub-workflows (fig. 1): an abstract simulation experiment containing processes generating data and metadata, metadata post-processing where heterogeneous metadata is processed into meaningful formats and structures, and usage of enriched data. Although the implementation of these processes can and will vary across different use cases, this generic knowledge production workflow encompasses archetypes of the components occurring in many implemented workflows.
@@
Figure 1: Metadata management in a generic knowledge production workflow. We conceptually divide the workflow into three consecutive sub-workflows: simulation, metadata post-processing, and data usage. Simulation collects raw metadata (yellow) at each step (blue), then stores it according to a data model describing the structures of data and raw metadata. Metadata post-processing (red) selects raw metadata, then parses and structures it; annotation links this processed metadata to the data. During post-processing, user-defined configuration specifies which raw metadata to select and how to process it, ensuring the final annotated data is suitable for its application. Some of these steps (light red) can be performed by the post-processing framework presented in “Metadata post-processing framework – the Archivist”. Data usage (green) analyzes data to draw conclusions through presentation and comparison; data may also be shared with others. Ultimately, the results of data usage inform subsequent simulations (green arrow). Each sub-workflow consists of autonomous steps (rectangles), user-configured steps (rounded rectangles), data storage (cylinders), and data transfer (arrows). This diagram aims to abstract simulation-based research; specific workflows may omit some elements (asterisks) or have additional steps.
@@
An abstract simulation experiment comprises a sequence of steps (fig. 1, Simulation section). We define these steps by summarizing the processes performed in two domain-specific workflows from neuroscience and hydrology[14, 20]. We consider the necessary steps to be configuration of the software environment, simulation engine, and model; execution of the simulation; and storage of the generated data and metadata; depending on the complexity of the preparation required before or after the execution of the simulation, there might be additional steps. Note that we do not claim that this abstract simulation experiment exhaustively represents all workflows or use cases.
@@
After a simulation experiment, users can analyze the resulting data and metadata to gain knowledge of the simulated model or simulation technology. The methods through which data is analyzed are tightly related to the objectives and intentions of the user. Furthermore, in some disciplines, these methods may require a complex workflow and perhaps even the use of HPC resources. Since we cannot exhaustively define the different existing strategies for processing data, we provide an abstract description of how the generated data can be exploited (fig. 1, Data usage section). For instance, data can be analyzed in order to measure metrics related to model predictions or accuracy, or simulation performance. Graphical visualizations are generated for easier human interpretation. Additionally, differences between multiple versions of simulators, models or parameters can be obtained by comparing with previous results. Then, newly obtained insights can be shared with peers to disseminate knowledge.
@@
Because processing the data may inadvertently remove or add specific features that alter the original information, storing the original data together with a data model allows independent users to exploit individual datasets with different methods, or to process multiple datasets at once, which increases re-usability of the simulation results. Over time, the data and metadata generated by different simulation experiments can be compiled and stored in a continuously growing collection. Establishing a data model is necessary to maintain the organization and integrity of this collection. A data model is an abstract representation of data objects and their relationships that organizes relevant attributes of data and standardizes how they relate to each other[21]. This serves as a blueprint for structuring data elements and their relationships, providing a clear understanding of how data is organized, stored, and utilized within a collection. Without the data model, users may not be able to establish an overview of the actual contents and may fail to efficiently explore the collection. For clarity we separate data and metadata storage in fig. 1, however this is an implementation choice left to the discretion of the user and for the remainder of this paper we refer to both storages together as one collection for simplicity.
@@
To implement and perform the diversity of data usage tasks, metadata is required to properly describe the simulation experiment. When exploring multiple sets of generated data, metadata categorizing each dataset allows researchers to efficiently differentiate them. Moreover, when sharing the results, metadata enhances the knowledge transfer between data users by detailing what, how, and where the experiment was performed. Metadata can exist at any step of the simulation (fig. 1, blue Data generation boxes), and can often be accessed only when performing the experiment. The collection of metadata and their interpretation often varies and evolves with the systems and software used (fig. 1, yellow Raw metadata collection boxes). We describe below some general practices for successfully collecting and handling the variety of metadata that arises during simulation (fig. 1, Metadata post-processing section).
@@
Metadata are data that describe data. Here, we focus on information that can be directly collected in the environment where the experiment takes place (sometimes termed "hard" metadata [22]). However, the definition of what precise pieces of information should be represented as metadata is difficult to generalize. Each experiment generates an arbitrary amount of information – separating this information into metadata and data depends on the nature and goal of the workflow itself. For instance, if the goal of a simulation workflow is to collect and analyze the data generated by the underlying model, then the performance of the simulation (time to solution, memory consumption, etc.) would be part of the metadata. On the contrary, if the goal of the workflow is to benchmark the simulator then the data generated by the model is of secondary importance, as the recorded performance data would be the desired primary result. Limiting the amount of information collected during a simulation experiment restricts how its output can be used in the future. Therefore, to maximize the potential for reuse of shared experiment results, it is important to collect as much information as possible[10]. Yet, the full extent of this collection needs to be determined on a case by case basis. Sometimes, collecting the full breadth of information available can be redundant. For example, when simulation models grow complex, the number of available parameters grows in volume too. Well-documented parameter sets often use default values suitable for their corresponding model; during experiments, only a subset of these may actually be modified. In this case, collecting only the modified parameters may be sufficient. Conversely, there may be instances where keeping detailed track of the differences between experiment is necessary. Sometimes, differences in software versions lead to discrepancies in experiment results[1]. In the case of rapidly evolving software such as the model, simulator, or their library dependencies, it is beneficial to keep track of the different versions used for each experiment. As a compromise between efficiency and completeness, we propose collecting information as sustainably as possible to an extent where all currently known user needs are satisfied.
@@
Given the rapidly evolving software and hardware platforms available for simulation alongside user-specific set-ups, it is impossible to exhaustively list all collectible metadata. Instead, we categorize the different sources of metadata according to each step of the simulation experiment (fig. 1, blue Data generation boxes):
@@
  • Software preparation: In this step all the required software (libraries, tools, and programs) are fetched, compiled if needed, and loaded into the environment. Information to collect here includes the source of the software, the parameters related to compilation, and the environment variables defined when loading the software.
  • Job generation : In HPC scenarios, job schedulers are generally used to efficiently allocate and share resources of compute clusters. Collectable metadata here include job parameters like machine name and partition used, amount and type of resources used, resource configuration such as process binding or thread pinning, and job time limit.
  • Pre-processing: Some simulation experiments require input data to calibrate or parametrize the model implementation. Using this input data may involve additional calculations or transformations within the workflow. If pre-processing is necessary, documenting the input data and procedures performed on it must be recorded to accurately determine the starting point of the simulation.
  • Job execution: The execution is the main part of the simulation, and involves instantiating the simulation model and running the simulation. It is at this step that recording the status of the execution environment is the most informative and least redundant. In contrast to Software preparation, the metadata collected in this step should focus on the environment variables, libraries, and processes present during execution. Additionally, performance information, resource consumption (memory, storage, network, etc.), and system load can only be recorded at this step.
  • Post-processing : Depending on the goal and implementation of the workflow, the data that is generated must be post-processed for further use, which may involve steps such as checking, cleaning, compressing, or transforming the data. If post-processing is necessary, documenting the procedures performed on data is essential to have an overview of how the final result was obtained. Thus information about expected output and the different methods used during this step must be recorded.
  • Data persistence: In this step, information on the storing process can be recorded. For example: who generated the data and metadata to be stored? When was it generated and when is it being stored? Where was it generated and where is it being stored e.g. a file server, a local or remote database? In the case of a file server, the address of the server and the absolute path of the stored files should be recorded. When using a database, new entries are indexed with a unique identifier; storing this identifier allows linking with future related entries and should be recorded correspondingly.
@@
There are two possible ways of collecting metadata: retrieval by a dedicated parallel process “pulling” the relevant pieces of information from each simulation step, or autonomous “pushing” by each component to a metadata storage. The pull-based approach relies on an overarching metadata-monitoring module that is adapted to the specific simulation. The push-based approach, in contrast, does not require such infrastructure at run time. At each step, the governing process simply dumps its metadata in some arbitrary, more or less raw format, independent of other simulation steps. The push-based approach is therefore much more flexible and can readily be applied to existing simulations. With each component presenting its own metadata it is likely that the final collection of metadata contains different structures and formats. Although it is possible to standardize the presentation of metadata across all components, this process would require transformations that are dependent on the objective for which the data is being generated, and can vary from user to user, or the interoperability standard that the researcher must adhere to. In some situations, transformations of metadata may even depend on other metadata, making it difficult to perform an ad-hoc transformation process. We therefore propose extracting and structuring the raw metadata in a separate step after the simulation. This presents a clear advantage by decoupling metadata collection processes from user goals and processing software. Collecting as much metadata as possible is essential as it is impossible to get the corresponding metadata after a simulation is performed. Thus, gathering the metadata regardless of format and storing it during simulation for later processing is an efficient way to keep track of the information describing the experiment. We call the collection of these unprocessed metadata files the “raw" metadata (fig. 1, yellow Raw metadata collection boxes).
@@
Transforming the heterogeneous raw metadata into a unified comprehensive format is essential in order to be able to exploit it alongside the data. Given that each workflow is run with a specific objective in mind, the collection of raw metadata often contains much more than needed. Although this is desirable to maximize possible reusability of simulation data, for a given objective the raw metadata should be filtered. For this, a dedicated read operation that extracts the relevant entries from a collection of metadata is performed according to the given objective. The extracted metadata is parsed and might be immediately usable depending on the output format of the parsing operations, however a specific structure might be needed to increase interoperability between the processed metadata and other data exploitation software. Therefore, structuring the metadata according to a standard is often a beneficial step to easily and unambiguously leverage the metadata information. With this, the previously heterogeneous metadata collection is now unified, possibly structured according to a standard, and ready to be exploited.
@@
Before being able to use processed metadata with data need to be combined . Here, we refer to combining as a referencing procedure in which any data entity that makes up a complete data record is cross-referenced with any metadata (item or collection) that assists in using and interpreting the data entity. Although it is possible to exploit metadata and data independently for specific purposes, combining them quickly becomes necessary when the volume and dimension of a dataset start to grow. When amassing data from multiple simulations, each with varying degrees of different configuration and/or models, it becomes apparent that simulation results need to be described at a higher level . For this, “tagging" or annotating the data with the processed metadata creates an enriched dataset. When appropriately stored, users can exploit indexing engines to navigate through enriched datasets and generate complex data projections. Again, the level of granularity with which data entities are identified and matched to metadata will need to be decided based on the objective of the workflow.
@@
Following the above practices will help in collecting and handling large heterogeneous volumes of metadata; additionally, combining datasets with processed metadata increases their explorability and exploitability. The actual methods through which these practices can be followed or implemented depend on the technology used to manage the simulation experiment and data processing pipeline. In particular some existing technologies offer support for these practices and are addressed in the “Discussion”.
Metadata post-processing framework – the Archivist
Users who follow the metadata collection practices outlined in the previous section might end up with large volumes of heterogeneous metadata. For example, users could potentially collect: code; documentation; system information; human readable and non-human readable data; binary, encoded, and compressed files; etc… Immediately exploiting the information found in this collection might be practically unfeasible due to the variety of file structure and formats. In particular, database systems and analysis tools often require specific input structures or formats without much flexibility. To be able to use their collected metadata, users need to individually transform the information in each file and then combine it into a specific output, or look for available software to aid in this process.
@@
Although there exist tools capable of doing these kinds of operations, e.g. AiiDA[15] parsers (https://.readthedocs.io/projects/-core/en/latest/reference/apidoc/.parsers.html, last access: 23 February 2024), and DataLad[17] metadata extractors (https://docs.datalad.org/projects/metalad/en/latest/extractors.html, last access: 23 February 2024), they may constrain how metadata is processed, and/or may be an inseparable component of a larger framework without the ability to be used independently.
Here, we propose the Archivist as a framework for performing parsing and structuring operations. Our goal is to facilitate the implementation of our metadata handling practices by providing the Archivist as an intermediary tool to extract information from a heterogeneous metadata file collection and unify it into a single file. Some users will want to structure the unified metadata file according to a vocabulary or schema such as RO-Crate[18], CodeMeta, or Bioschemas[19]. Whereas our proposed metadata practices aim to collect information describing the conditions under which an individual result was obtained, vocabularies and schemas provide a higher level of description for entities, context, and goals surrounding the experiment itself. Nonetheless, some of the collected metadata can help to generate documents according to these defined standards. To do this, users can define data templates to structure the extracted information from the collection of metadata files into a comprehensive output compliant with their desired standard. Still, users will likely need to provide additional input to fully exploit the expressivity of the vocabularies. An example output compliant with the Bioschemas vocabulary for computational workflows is shown in “Methods”.
@@
The Archivist doesn’t currently support data annotation, i.e., linking metadata records to relevant parts of the data in the post-processing pipeline (fig. 1, red Annotation box), because the heavy customizations required to fit the underlying structure of the data are difficult to formalize. For example, while data structured according to HDF5-based[23] formats like the NIX[24] standard can make use of built-in mechanisms for linking collections of metadata entries to data, such mechanisms cannot necessarily be generalized to other file formats and data models.